Is this correct
solving for x


5^3x= 5^(4x-2)
x- (log⁡(25)+2 i π n)/(log⁡(5))
x=2



Answer by Alan3354(69443)

Is this correct
solving for x


5^3x= 5^(4x-2)
x- (log⁡(25)+2 i π n)/(log⁡(5))
x=2
==============
I'm not sure what you did.
5^3x= 5^(4x-2)
3x = 4x - 2
x = 2